1. Solve the given differential equation by undetermined coefficients.
y'' − 2y' + 5y = e^x sin x

y(x) =____

2. Solve the given differential equation by undetermined coefficients.
y'' − y' − 2y = e^2x

y(x) =___
1
Expert's answer
2020-06-30T18:01:02-0400

1.The given equation is "(D^{2}-2D+5)y = e^{x}\sin x".

The auxiliary equation is "m^{2}-2m+5=0." Solving for m, we get "m=1\pm 2i".

Thus, the complementary function is "C.F = e^{x}(c_{1}\cos 2x + c_{2}\sin 2x)."


To find the particular integral, let "y_{p} = Ae^{x}\sin x" be the trial solution.

Then,

"y'_{p} = Ae^{x}(\cos x + \sin x)~\text{and}~\\ 
y''_{p} = A\bigg(e^{x}(-\sin x+\cos x )+e^{x}(\cos x + \sin x)\bigg)\\ ~~~~~= 2Ae^{x}\cos x"


Therefore,

"2Ae^{x}\cos x - 2\bigg(Ae^{x}(\cos x+\sin x)\bigg)+5Ae^{x}\sin x = e^{x}\sin x\\
A = \dfrac{1}{3}"

Hence the particular integral is "y_{p} = \dfrac{1}{3}e^{x}\sin x" .

The general solution is "y=e^{x}(c_{1}\cos 2x+c_{2}\sin 2x)+\dfrac{1}{3}e^{x}\sin x"


2.The given equation is "(D^{2}-D-2)y=e^{2x}."

The auxiliary equation is "m^{2}-m-2=0". Solving for m, we get "m=-1,2".

Thus, the complementary function is "C.F=c_{1}e^{-x}+c_{2}e^{2x}".


To find the particular integral, since "g(x) =e^{2x}" is a linear combination of a term in complementary function we choose "y_{p} = Axe^{2x}" as the trial particular solution.

Thus,

"y'_{p}=A(2xe^{x}+e^{2x})=Ae^{2x}(2x+1)\\
y''_{p}=A\bigg(2(2xe^{x}+e^{2x})+2e^{2x}\bigg)=4Ae^{2x}(x+1)\\"


Therefore,

"4Ae^{2x}(x+1)-Ae^{2x}(2x+1)-2Axe^{2x}=e^{2x}\\
A(4x+4-2x-1-2x)=1\\
A=\dfrac{1}{3}"

Hence the particular integral is, "y_{p}=\dfrac{1}{3}xe^{2x}"


The general solution is, "y=c_{1}e^{-x}+e^{2x}(\frac{x}{3}+c_{2})"